OpenStudy (anonymous):

find the square of this binomial [(xy+9)-5]^2

jimthompson5910 (jim_thompson5910):

[(xy+9)-5]^2 = [(xy+9)-5][(xy+9)-5] [(xy+9)-5]^2 = (xy+9)^2-5(xy+9)-5(xy+9)+25 [(xy+9)-5]^2 = (xy+9)^2-10(xy+9)+25 [(xy+9)-5]^2 = (xy)^2+18xy+81-10xy-90+25 [(xy+9)-5]^2 = x^2y^2+8xy+16

OpenStudy (anonymous):

thanks so much!!

jimthompson5910 (jim_thompson5910):

np